5. What force is needed to pull one microscope slide over another at a speed of 1.00 cm/s, if there is a 0.500-mm-thick layer of 20 degrees C water between them and the contact area is 8.50 cm^2?
Answer #1

Solve Given V = 1.00(m/s = 0.01 m/s L=0.500mm L=0.500Xlom A = 8.500m² = 8.50 x 10 m2 Viscosity of wader at 20°C 2 = 1 mpa.s =